OpenStudy (anonymous):

how do i solve the equation: (x-5)(x+7)=0

OpenStudy (anonymous):

think which x values would make it true

OpenStudy (anonymous):

-5 i believe i am not sure

OpenStudy (jim766):

(x-5)(x + 7) = 0 the zero product property says that if 2 things are mult and the answer is zero then one or both things must be = 0 so we can do this x - 5 = 0 or x + 7 = 0 now just solve each for x

OpenStudy (jim766):

do each separately, tell me what you get

OpenStudy (anonymous):

X=5 and the other X=-7

OpenStudy (jim766):

great!

OpenStudy (anonymous):

thank you for your help

OpenStudy (jim766):

yw